# Terrain Diffusion: A Diffusion-Based Successor to Perlin Noise in Infinite, Real-Time Terrain Generation

### v2

@@ A.2. Additional Details on FID Calculations
For Perlin blending (Jain et al., 2023), our algorithm is not identical to the one used in the original source, since the original source targets a different, zero-centered dataset. For fair comparison, we adapt the method for our dataset. While the original algorithm applies Perlin noise with a static distribution, we utilize an adaptive distribution that is centered around the mean of the surrounding elevation tiles, and scaled by the standard deviation of these tiles. We utilize linear interpolation to smoothly interpolate between the distributions of neighboring tiles.
@@ B.1. Preliminaries and Notation
Algorithm 1 efficiently computes (3). At level $t$ it maintains images $A_{t-1},B_{t-1}\in\mathcal{J}$ and a processed set $P_{t-1}\subseteq S$. To answer a query $J_{t-1}[R]$, it performs:
For each window $i\in\kappa(R)\setminus P_{t-1}$:
$A_{t-1}[R_{i}]$ $\leftarrow A_{t-1}[R_{i}]+W_{i}\otimes\Phi(J_{t}[R_{i}],y_{i}),$
$B_{t-1}[R_{i}]$ $\leftarrow B_{t-1}[R_{i}]+W_{i}.$
Add each such $i$ to $P_{t-1}$. The result is
$J_{t-1}[R]=A_{t-1}[R]/B_{t-1}[R].$
Where the division is performed elementwise. New windows are evaluated recursively by querying $J_{t}[\cdot]$ in the same way at level $t$, until reaching $J_{T}^{(s)}$, which is given by the seed. At intialization, $A_{t-1}=\textbf{0}$, $B_{t-1}=\textbf{0}$, and $P_{t-1}=\emptyset$.
@@ B.2. Seed Consistency
Fix $s\in\mathcal{X}$ and a timestep $t\in\{1,\dots,T\}$. Then immediately before any query $J_{t-1}[R]$ with Algorithm 1, the pair $(A_{t-1},B_{t-1})$ satisfy
$A_{t-1}$ $=\sum_{i\in P_{t-1}}U_{i}\bigl(W_{i}\otimes\Phi(J_{t}^{(s)}[R_{i}]\mid y_{i}^{(s)})\bigr),$
$B_{t-1}$ $=\sum_{i\in P_{t-1}}U_{i}(W_{i}).$
Furthermore, after performing one iteration of Algorithm 1, the updated state satisfies the same form with $P_{t-1}$ replaced by $P_{t-1}\cup\kappa(R)$.
@@
We prove this by induction over an arbitrary sequence of queries. At initialization, $A_{t-1}=0$, $B_{t-1}=0$, and $P_{t-1}=\emptyset$, so the claim is trivially true.
@@
During the query, for each $i\in\kappa(R)\setminus P_{t-1}$ the algorithm performs
$A_{t-1}[R_{i}]\leftarrow A_{t-1}[R_{i}]+V_{i},\qquad B_{t-1}[R_{i}]\leftarrow B_{t-1}[R_{i}]+W_{i}$
@@
$A_{t-1}\leftarrow A_{t-1}+U_{i}(V_{i}),\qquad B_{t-1}\leftarrow B_{t-1}+U_{i}(W_{i}).$
@@
$A_{t-1}=\sum_{i\in P_{t-1}}U_{i}(V_{i})+\sum_{i\in\kappa(R)\setminus P_{t-1}}U_{i}(V_{i})=\sum_{i\in P_{t-1}\cup\kappa(R)}U_{i}(V_{i}),$
@@
$B_{t-1}=\sum_{i\in P_{t-1}\cup\kappa(R)}U_{i}(W_{i}).$
Finally, the algorithm updates $P_{t-1}\leftarrow P_{t-1}\cup\kappa(R)$, so the updated state satisfies the same form with $P_{t-1}$ replaced by $P_{t-1}\cup\kappa(R)$.
By induction, the claim holds for all queries $J_{t-1}[R]$. ∎
@@
$A_{t-1}=\sum_{i\in P^{\prime}_{t-1}}U_{i}(V_{i}),\qquad B_{t-1}=\sum_{i\in P^{\prime}_{t-1}}U_{i}(W_{i}),$
for some processed set $P^{\prime}_{t-1}\supseteq\kappa(R)$.
@@
$A_{t-1}[R]=\left(\sum_{i\in P^{\prime}_{t-1}}U_{i}(V_{i})\right)[R]=\left(\sum_{i\in\kappa(R)}U_{i}(V_{i})\right)[R],$
@@
$B_{t-1}[R]=\left(\sum_{i\in\kappa(R)}U_{i}(W_{i})\right)[R].$
@@ B.3. Constant-Time Random Access
For any window index $i$, Theorem B.7 gives a uniform bound $C_{0}(R_{i})\leq K$ on the number of calls to $\Phi$ needed to answer a query for $J_{0}[R_{i}]$, starting from an empty cache. This bound depends only on $T$ and $M$, which are constants, and not on $i$ or any regions previously processed. In standard algorithmic notation, this means that the time complexity of a query $J_{0}[R]$ is $O(1)$, when $R$ is a window region.
When caches are reused across multiple queries, subsequent queries typically cost far less than $K$, but this is not required for the asymptotic guarantee.
@@ B.4. Parallelization
Recall that, at fixed $t$ and seed $s$, the numerator and denominator images for the update $J_{t}^{(s)}\mapsto J_{t-1}^{(s)}$ are
$A_{t-1}^{(s)}$ $=\sum_{i\in S}U_{i}\bigl(W_{i}\otimes\Phi(J_{t}^{(s)}[R_{i}]\mid y_{i}^{(s)})\bigr),$
$B_{t-1}^{(s)}$ $=\sum_{i\in S}U_{i}(W_{i}),$
@@
$J_{t-1}^{(s)}=A_{t-1}^{(s)}/B_{t-1}^{(s)}.$
We treat calls to $\Phi$ as the only expensive operation, and all other operations (additions, multiplications, divisions and updates to $A_{t-1},B_{t-1}$) as free. A computation is parallelizable if there exists an algorithm in which all calls to $\Phi$ can be partitioned into finitely many rounds so that within each round the calls are independent and can be executed simultaneously.
We first note that, at a fixed timestep, once the inputs $J_{t}^{(s)}[R_{i}]$ are known, all required model evaluations can be done in parallel.
@@
Fix a seed $s$, a timestep $t\in\{1,\dots,T\}$, and a finite set of window indices $I\subseteq S$. Suppose that for every $i\in I$ the tensor $J_{t}^{(s)}[R_{i}]$ is available for free. Then the evaluations
$\Phi(J_{t}^{(s)}[R_{i}]\mid y_{i}^{(s)}),\qquad i\in I,$
can be performed in a single parallel round, and all corresponding contributions to $A_{t-1}^{(s)}$ and $B_{t-1}^{(s)}$ on $\bigcup_{i\in I}R_{i}$ can be formed without any further calls to $\Phi$.
@@
$A_{t-1}^{(s)}$ $\leftarrow A_{t-1}^{(s)}+U_{i}\bigl(W_{i}\otimes\Phi(J_{t}^{(s)}[R_{i}]\mid y_{i}^{(s)})\bigr)$
$B_{t-1}^{(s)}$ $\leftarrow B_{t-1}^{(s)}+U_{i}(W_{i})$
and the division $J_{t-1}^{(s)}=A_{t-1}^{(s)}/B_{t-1}^{(s)}$ involve only element-wise arithmetic and therefore require no additional model evaluations. Hence all work associated with the windows in $I$ can be completed using a single parallel round of calls to $\Phi$. ∎
@@
$\Phi(J_{t}^{(s)}[R_{i}]\mid y_{i}^{(s)}),\qquad i\in I_{t},$
required to construct $A_{t-1}^{(s)}$ and $B_{t-1}^{(s)}$ on $R^{\ast}$ can be carried out in a single additional parallel round. No more model evaluations are needed to extract $J_{t}^{(s)}[R^{(k)}]$ from these images.
@@ C.3. Dependency Chaining and DAGs
Appendix D Dataset Details
@@
Appendix E Signed Square-Root Transform

### v3

@@ A.2. Additional Details on FID Calculations
For Perlin blending (Jain et al., 2023), our implementation differs from the reference method, which targets a zero-centered dataset. To ensure a fair comparison, we adapt the algorithm to our data: rather than applying Perlin noise with a static distribution, we employ an adaptive distribution centered on the mean and scaled by the standard deviation of surrounding elevation tiles (weighted by linear interpolation).
@@ B.1. Preliminaries and Notation
Algorithm 1 efficiently computes (3). At level $t$ it maintains images $A_{t},B_{t}\in\mathcal{J}$ and a processed set $P_{t}\subseteq S$. To answer a query $J_{t}[R]$, it performs:
For each window $i\in\kappa(R)\setminus P_{t}$:
$A_{t}[R_{i}]$ $\leftarrow A_{t}[R_{i}]+W_{i}\otimes\Phi(J_{t}[R_{i}],y_{i}),$
$B_{t}[R_{i}]$ $\leftarrow B_{t}[R_{i}]+W_{i}.$
Add each such $i$ to $P_{t}$. The result is
$J_{t}[R]=A_{t}[R]/B_{t}[R].$
Where the division is performed elementwise. New windows are evaluated recursively by querying $J_{t}[\cdot]$ in the same way at level $t$, until reaching $J_{T}^{(s)}$, which is given by the seed. At intialization, $A_{t}=\textbf{0}$, $B_{t}=\textbf{0}$, and $P_{t}=\emptyset$.
@@ B.2. Seed Consistency
Fix $s\in\mathcal{X}$ and a timestep $t\in\{1,\dots,T\}$. Then immediately before any query $J_{t}[R]$ with Algorithm 1, the pair $(A_{t},B_{t})$ satisfy
$A_{t}$ $=\sum_{i\in P_{t}}U_{i}\bigl(W_{i}\otimes\Phi(J_{t}^{(s)}[R_{i}]\mid y_{i}^{(s)})\bigr),$
$B_{t}$ $=\sum_{i\in P_{t}}U_{i}(W_{i}).$
Furthermore, after performing one iteration of Algorithm  1, the updated state satisfies the same form with $P_{t}$ replaced by $P_{t}\cup\kappa(R)$.
@@
We prove this by induction over an arbitrary sequence of queries. At initialization, $A_{t}=0$, $B_{t}=0$, and $P_{t}=\emptyset$, so the claim is trivially true.
@@
During the query, for each $i\in\kappa(R)\setminus P_{t}$ the algorithm performs
$A_{t}[R_{i}]\leftarrow A_{t}[R_{i}]+V_{i},\qquad B_{t}[R_{i}]\leftarrow B_{t}[R_{i}]+W_{i}$
@@
$A_{t}\leftarrow A_{t}+U_{i}(V_{i}),\qquad B_{t}\leftarrow B_{t}+U_{i}(W_{i}).$
@@
$A_{t}=\sum_{i\in P_{t}}U_{i}(V_{i})+\sum_{i\in\kappa(R)\setminus P_{t}}U_{i}(V_{i})=\sum_{i\in P_{t}\cup\kappa(R)}U_{i}(V_{i}),$
@@
$B_{t}=\sum_{i\in P_{t}\cup\kappa(R)}U_{i}(W_{i}).$
Finally, the algorithm updates $P_{t}\leftarrow P_{t}\cup\kappa(R)$, so the updated state satisfies the same form with $P_{t}$ replaced by $P_{t}\cup\kappa(R)$.
By induction, the claim holds for all queries $J_{t}[R]$. ∎
@@
$A_{t}=\sum_{i\in P^{\prime}_{t}}U_{i}(V_{i}),\qquad B_{t}=\sum_{i\in P^{\prime}_{t}}U_{i}(W_{i}),$
for some processed set $P^{\prime}_{t}\supseteq\kappa(R)$.
@@
$A_{t}[R]=\left(\sum_{i\in P^{\prime}_{t}}U_{i}(V_{i})\right)[R]=\left(\sum_{i\in\kappa(R)}U_{i}(V_{i})\right)[R],$
@@
$B_{t}[R]=\left(\sum_{i\in\kappa(R)}U_{i}(W_{i})\right)[R].$
@@ B.3. Constant-Time Random Access
For any window index $i$, Theorem B.7 gives a uniform bound $C_{0}(R_{i})\leq K$ on the number of calls to $\Phi$ needed to answer a query for $J_{0}[R_{i}]$, starting from an empty cache. This bound depends only on $T$ and $M$, which are constants, and not on $i$ or any regions previously processed. In standard algorithmic notation, this means that the time complexity of a query $J_{0}[R]$ is $O(1)$, when $R$ is a window region. Queries typically cost far less than $K$, but this is not required for the asymptotic guarantee.
@@ B.4. Parallelization
Recall that, at fixed $t$ and seed $s$, the numerator and denominator images for the update $J_{t+1}^{(s)}\mapsto J_{t}^{(s)}$ are
$A_{t}^{(s)}$ $=\sum_{i\in S}U_{i}\bigl(W_{i}\otimes\Phi(J_{t+1}^{(s)}[R_{i}]\mid y_{i}^{(s)})\bigr),$
$B_{t}^{(s)}$ $=\sum_{i\in S}U_{i}(W_{i}),$
@@
$J_{t}^{(s)}=A_{t}^{(s)}/B_{t}^{(s)}.$
We treat calls to $\Phi$ as the only expensive operation, and all other operations (additions, multiplications, divisions and updates to $A_{t},B_{t}$) as free. In this work, a computation is parallelizable if there exists an algorithm in which all calls to $\Phi$ can be partitioned into a constant number of rounds so that within each round the calls are independent and can be executed simultaneously.
We first note that, at a fixed timestep, once the inputs $J_{t+1}^{(s)}[R_{i}]$ are known, all required model evaluations can be done in parallel.
@@
Fix a seed $s$, a timestep $t\in\{0,\dots,T-1\}$, and a finite set of window indices $I\subseteq S$. Suppose that for every $i\in I$ the tensor $J_{t+1}^{(s)}[R_{i}]$ is available for free. Then the evaluations
$\Phi(J_{t+1}^{(s)}[R_{i}]\mid y_{i}^{(s)}),\qquad i\in I,$
can be performed in a single parallel round, and all corresponding contributions to $A_{t}^{(s)}$ and $B_{t}^{(s)}$ on $\bigcup_{i\in I}R_{i}$ can be formed without any further calls to $\Phi$.
@@
$A_{t}^{(s)}$ $\leftarrow A_{t}^{(s)}+U_{i}\bigl(W_{i}\otimes\Phi(J_{t}^{(s)}[R_{i}]\mid y_{i}^{(s)})\bigr)$
$B_{t}^{(s)}$ $\leftarrow B_{t}^{(s)}+U_{i}(W_{i})$
and the division $J_{t}^{(s)}=A_{t}^{(s)}/B_{t}^{(s)}$ involve only element-wise arithmetic and therefore require no additional model evaluations. Hence all work associated with the windows in $I$ can be completed using a single parallel round of calls to $\Phi$. ∎
@@
$\Phi(J_{t+1}^{(s)}[R_{i}]\mid y_{i}^{(s)}),\qquad i\in I_{t},$
required to construct $A_{t}^{(s)}$ and $B_{t}^{(s)}$ on $R^{\ast}$ can be carried out in a single additional parallel round. No more model evaluations are needed to extract $J_{t}^{(s)}[R^{(k)}]$ from these images.
@@ C.3. Dependency Chaining and DAGs
Appendix D Additional Implementation Details
An overview of our architecture with hyperparameters is provided in Table  5.
The generative components of Terrain Diffusion are built upon the EDM2 architecture (Karras et al., 2024b). We modify the base architecture, following sCM  (Lu and Song, 2025), by implementing pixel-norm on embedding vectors and substituting Fourier embeddings with Positional embeddings. To incorporate auxiliary conditioning, we utilize a magnitude-preserving linear layer to project flattened conditioning inputs
Fidelity Optimization.
For both the core latent and decoder models, we optimize fidelity using AutoGuidance (Karras et al., 2024a). Hyperparameter sweeps are conducted via Optuna, minimizing Kernel Inception Distance (KID) (Bińkowski et al., 2018) calculated over 1,000 samples. We terminate the sweeps after 300 runs, or once KID improvement saturates.
Consistency Distillation.
The core latent model and decoder are distilled into continuous-time consistency models (sCM). We perform post-hoc sweeps for the EMA parameter ($\sigma_{\text{rel}}$) and the optimal stopping iteration. This prevents fidelity degradation observed during late-stage training.
Training Speed
The entire pipeline can be trained from scratch on a top-tier consumer-grade GPU (e.g., RTX 5090) in approximately one week, enabling practical retraining on custom datasets. If the decoder is reused, fine tuning the base model can be completed in about a day.
Table 5. Hyperparameters and architectural details for the component models.
Parameter Core Latent Consistency Decoder Coarse Model Autoencoder
Base Architecture EDM2 + sCM EDM2 + sCM EDM2 + sCM EDM2 + sCM - Skips
Model Channels 192 (128 Guide) 64 (32 Guide) 128 64
Channel Multipliers $[1,2,3,4]$ $[1,2,3,4]$ $[1,2,3,4]$ $[1,2,4,4]$
Layers per Block 3 3 2 2
Training Crop Size $64\times 64$ (Latent) $128\times 128$ $16\times 16$ $128\times 128$
Dropout 10% 0% 10% 0%
LR (Training) $5\times 10^{-3}$ ($1\times 10^{-2}$ Guide) $\times 10^{-2} 1×1$$\times 10^{-2}$ $1\times 10^{-2}$
LR (Distillation) $5\times 10^{-5}$ $1\times 10^{-4}$
Batch Size 48 (64 Guide) 64 256 64
Training Steps 1,024,000 716,800 153,600 153,600
Guide Training Steps 102,400 102,400
Distillation Steps 25,600 25,600
Guidance AutoGuidance AutoGuidance
Guidance Hyperparameters
Main EMA $\sigma_{\text{rel}}$ 0.12 0.093
Guide EMA $\sigma_{\text{rel}}$ 0.14 0.243
Guide Stopping Step 57,344 70,656
Guidance Scale 2.15 1.26
Consistency / Distillation Hyperparameters
Intermediate Timestep ($\sigma$) 0.35 One-Step or 0.065
$\sigma_{\text{rel}}$ 0.08 0.21
EMA Step 69,632 60,416
Appendix E Dataset Details
@@
Appendix F Signed Square-Root Transform